OpenStudy (anonymous):

sqrt128+2sqrt112-2sqrt98-628 solve pweaseeee

OpenStudy (anonymous):

8√2 +2x4 √7 -2x7√2 -628 -6√2 + 8√7 -628 is that 628 use sqrt too?

OpenStudy (anonymous):

no 628 isnt sqrt too , thank you :)

OpenStudy (anonymous):

your welcome :D

OpenStudy (anonymous):

\[\sqrt{2*64}+2\sqrt{8*14}-2\sqrt{2*49}-628\] \[8\sqrt{2}+8\sqrt{14}-14\sqrt{2}-628\] \[8\sqrt{14}-6\sqrt{2}-628\]

OpenStudy (anonymous):

just kiddig i just looked back at the problem and it is actually 6sqrt28.. im orry for my mistake

OpenStudy (anonymous):

just kidding!!!!! is this question is for fun?????????

OpenStudy (anonymous):

if that so 6 √28 = 12√7 so = -6√2+20√7

OpenStudy (anonymous):

thank youuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuu your a life saver!

OpenStudy (anonymous):

nope he/she is a 'star'